OpenStudy (luigi0210):

Solve the differential equation:

OpenStudy (luigi0210):

\[\frac{ d^2y }{ dx^2 }+2\frac{ dy }{ dx }+2y=10\]

OpenStudy (anonymous):

Laplace transforms

OpenStudy (jhannybean):

\[\large y'' +y'+2y = 10\] how do you do laplace transformations..... :|

OpenStudy (anonymous):

Wait NVM LOL Do you have any initial conditions?

OpenStudy (jhannybean):

YEAH got you there. :|

OpenStudy (luigi0210):

what's that?

OpenStudy (jhannybean):

what's what?

OpenStudy (luigi0210):

Laplace transformations..

OpenStudy (jhannybean):

Idk....

OpenStudy (luigi0210):

or laplace transforms..

OpenStudy (reemii):

\(y''++y'+2y=10\) 1) Homogeneous equation: \(y''+y+2y=0\) (H) -> \(p(\lambda) = \lambda^2 + \lambda + 2=0\) iff \(\lambda=\frac{-1\pm i\sqrt7}{2} = -\frac12 \pm i\frac{\sqrt7}2\) (i think) The functions \(g(x) = e^{-\frac12x}(A...+B...)\), with \(A,B\in\mathbb R\) are solutions of (H). 2) find a particular solution. It seems that one constant function is solution of this. (\(y(x) = c\) for some \(c\)). 3) write the general form of the solution. This method is quite common.

OpenStudy (jhannybean):

@reemii that looks way too complicated...

OpenStudy (reemii):

it's a method, no need to think. It is likely the asker learned the method. ( @Luigi0210 )

OpenStudy (luigi0210):

hm, from what math is it?

OpenStudy (reemii):

what do you mean?

OpenStudy (luigi0210):

Is this method from calculus 2?

OpenStudy (reemii):

no idea.

OpenStudy (reemii):

@Loser66 from \(\lambda=-\frac12\pm i \sqrt{7}/2\) you write \(u=-\frac12,v=\sqrt{7}/2\). then (complete version) \(g(x)=e^{ux}(A\cos(vx)+B\sin(vx))\). A solution of the initial equation is \(y(x)=5\). General solution: \(y_\text{gen}(x)=5+g(x)\), \(A,B\in\mathbb R\).

OpenStudy (reemii):

I'll look into the laplace transorm method. :)

OpenStudy (anonymous):

Can't use laplace transforms here without initial conditions. Sry :(

OpenStudy (anonymous):

@Loser66 Convert your 2nd order ODE into two 1st order ODE's

OpenStudy (loser66):

@FutureMathProfessor the same with my method.need initial conditions

OpenStudy (loser66):

@FutureMathProfessor You know that I didn't take differential equation yet, how can I know what is Laplace transform? I use my own knowledge only.

OpenStudy (anonymous):

@Loser66 This is my DiffEQ test from a VERY similar problem. No initial conditions needed http://gyazo.com/e8615a0188dd35317e433142262c032c.png